# AutoDroid-V2: Boosting SLM-based GUI Agents via Code Generation

### v2

@@ 3.2.3. Tree-based Script Quality Improvement
Even if a script executes successfully, it may not align with the intended user tasks. Thus, AutoDroid-V2 uses a on-cloud LLM-based reward model to evaluate whether the code $C$ successfully achieves the task $\mathcal{T}$. This reward model takes as input the task $\mathcal{T}$, the executed code $C$, and a sequence of environment observation states $\mathcal{S}=(s_{1},\dots,s_{n})$, outputting a judgment $\mathcal{J}$ and feedback $\mathcal{F}_{\text{out}}$. If the reward model deems the code $C$ as complete, the task-code pair $(\mathcal{T},C)$ is added to the dataset. Otherwise, the error information $\mathcal{E}$ and feedback $\mathcal{F}_{\text{out}}$ are recorded and used in a code search tree to regenerate a code sample.
@@ 4. Evaluation
We implement AutoDroid-V2 using Python and Java. SLMs are deployed on mobile devices using llama.cpp, a framework to enable LLM/SLM inference in C/C++.
@@ 4.1. Experimental Setup
Dataset. We evaluate the effectiveness and efficiency of AutoDroid-V2 and baseline on two benchmarks: DroidTask (Wen et al., 2024) and AitW-subset (Rawles et al., 2023). DroidTask (Wen et al., 2024) is a mobile task automation dataset with 158 high-level tasks across 13 popular apps. AitW (Zhang et al., 2024b) is a large-scale dataset for Android device control. The original AitW dataset contains 417k tasks but lacks the exploration environment of the apps. So we choose the ‘google apps’ and ‘general’ subsets of the AitW dataset with 68 tasks using LlamaTouch (Zhang et al., 2024b) environment, which is executable and can be explored. The exploration trace is provided by the original datasets. Our document generation and task synthesis are independent of the test set.
@@
  • Success Rate: In DroidTask (Wen et al., 2024), a task is considered completed if the ground-truth action sequence $\mathcal{A}$ is a subsequence of the agent-generated sequence $\hat{\mathcal{A}}$, i.e., $\mathcal{A}\in\hat{\mathcal{A}}$. This metric reflects the agent’s ability to complete the task. The LlamaTouch framework (Zhang et al., 2024b) evaluates task completion by sequentially comparing a task execution trace $\{\hat{U}_{1},\hat{U}_{2},...,\hat{U}_{k}\}$ against selected essential states $\{U_{i_{1}},U_{i_{2}},...\}$. A task is considered complete if the task execution trace includes all the essential states in order.
@@ 4.3. Latency and Cost
Table 5 shows the average runtime token consumption for AutoDroid-V2 and the baseline method. The cached tokens are the common prefix part that has already been processed, and the internal state of LLM has been stored in a file, which can be pre-loaded to memory and reused. The cached tokens refer to the common prefix that has already been processed, with the LLM’s internal state stored in a file, allowing it to be preloaded into memory and reused. The remaining tokens are part of the input prompt following the cached prefix, which must be processed in real time during model inference. For step-wise method, the main part of a prompt is the GUI state, which cannot be cached because they are dynamic and frequent changes. Conversely, AutoDroid-V2 primarily uses the app document, which is static and can be cached. The latency for each response mainly depends on the remaining and output tokens. AutoDroid-V2 saves 97.8% of the remaining tokens and 85.2% of the output tokens, significantly reducing the computational burden.
@@ 4.4.1. Performance with Different LLMs
We use AutoDroid, the best-performing baseline on the Llama-3.8B model, as a reference and compare it with AutoDroid-V2 across different LLMs. The results presented in Figure 5 show the comparison of success rate and reversed redundancy ratio between AutoDroid-V2 and the baseline models for different LLMs: Llama3.2-3B, Qwen2.5-7B, and Llama3.1-8B, where AutoDroid-V2 achieves 44.6%, 50.0%, 54.4% success rate respectively. The reversed redundancy ratio of AutoDroid-V2 is stable, ranging from 90.5% to 93.0%. This is due to the generated scripts are designed to be straightforward and focused on task completion without including exploration or trial-and-error steps.
@@ 5. Discussion
Another possible limitation of AutoDroid-V2 is the difficulty to handle highly dynamic user interfaces, such as those in web search engines and games. In these apps, it is difficult to plan the actions in advance, since the future states are usually unpredictable. In concept, AutoDroid-V2 can deal with such dynamic user interfaces, but it may need to regenerate the plan for each step, leading to degraded efficiency. Our future work will focus on how to seamlessly integrate the step-wise reasoning approach with our script-based approach. Future agents are expected to flexibly switch between the two modes for different types of user interfaces, handling dynamic content effectively while maintaining the efficiency.
@@ 6. Conclusion
7. ACKNOWLEDGEMENT
This work is supported by National Natural Science Foundation of China (Grant No.62272261), the National Key R&D Program of China (No.2022YFF0604501), Xiaomi Foundation, and Tsinghua University (AIR)–AsiaInfo Technologies (China) Inc. Joint Research Center.

### v3

@@ 3.2.3. Tree-based Script Quality Improvement
Even if a script executes successfully, it may not align with the intended user tasks. Thus, AutoDroid-V2 uses an on-cloud LLM-based reward model (Zheng et al., 2023) to evaluate whether the code $C$ successfully achieves the task $\mathcal{T}$. This reward model takes as input the task $\mathcal{T}$, the executed code $C$, and a sequence of environment observation states $\mathcal{S}=(s_{1},\dots,s_{n})$, outputting a judgment $\mathcal{J}$ and feedback $\mathcal{F}_{\text{out}}$. If the reward model deems the code $C$ as complete, the task-code pair $(\mathcal{T},C)$ is added to the dataset. Otherwise, the error information $\mathcal{E}$ and feedback $\mathcal{F}_{\text{out}}$ are recorded and used in a code search tree to regenerate a code sample.
@@ 4. Evaluation
We implement AutoDroid-V2 using Python and Java. SLMs are deployed on mobile devices using llama.cpp, a framework to enable LLM/SLM inference in C/C++. AutoDroid-V2 collects current GUI information via ADB (Android Debug Bridge). For each GUI action API, ADB commands are sent to the identified GUI element.
@@ 4.1. Experimental Setup
Dataset. We evaluate the effectiveness and efficiency of AutoDroid-V2 and baseline on two benchmarks: DroidTask (Wen et al., 2024) and AitW-subset (Rawles et al., 2023). DroidTask (Wen et al., 2024) is a mobile task automation dataset with 158 high-level tasks across 13 popular apps. AitW (Zhang et al., 2024c) is a large-scale dataset for Android device control. The original AitW dataset contains 417k tasks but lacks the exploration environment of the apps. So we choose the ‘google apps’ and ‘general’ subsets of the AitW dataset with 68 tasks using LlamaTouch (Zhang et al., 2024c) environment, which is executable and can be explored. The exploration trace is provided by the original datasets. Our document generation and task synthesis are independent of the test set.
Training Details. We perform supervised fine-tuning on our model using the automatically generated task-solution training set detailed in Section 3.2, training for one epoch with a batch size of 32, a learning rate of 1e-5, and a maximum sequence length of 8192. For the baseline AutoDroid (Wen et al., 2024), we use the same training parameters.
@@
  • Success Rate: In DroidTask (Wen et al., 2024), a task is considered completed if the ground-truth action sequence $\mathcal{A}$ is a subsequence of the agent-generated sequence $\hat{\mathcal{A}}$, i.e., $\mathcal{A}\in\hat{\mathcal{A}}$. This metric reflects the agent’s ability to complete the task. The LlamaTouch framework (Zhang et al., 2024c) evaluates task completion by sequentially comparing a task execution trace $\{\hat{U}_{1},\hat{U}_{2},...,\hat{U}_{k}\}$ against selected essential states $\{U_{i_{1}},U_{i_{2}},...\}$. A task is considered complete if the task execution trace includes all the essential states in order.
@@ 4.3. Latency and Cost
Table 4 shows the average runtime token consumption for AutoDroid-V2 and the baseline method. The cached tokens are the common prefix part that has already been processed, and the internal state of LLM has been stored in a file, which can be pre-loaded to memory and reused. The cached tokens refer to the common prefix that has already been processed, with the LLM’s internal state stored in a file, allowing it to be preloaded into memory and reused. The remaining tokens are part of the input prompt following the cached prefix, which must be processed in real-time during model inference. For the step-wise method, the main part of a prompt is the GUI state, which cannot be cached because they are dynamic and frequent changes. Conversely, AutoDroid-V2 primarily uses the app document, which is static and can be cached. The latency for each response mainly depends on the remaining and output tokens. AutoDroid-V2 saves 97.8% of the remaining tokens and 85.2% of the output tokens, significantly reducing the computational burden.
@@ 4.4.1. Performance with Different LLMs
We use AutoDroid, the best-performing baseline on the Llama-3.8B model, as a reference and compare it with AutoDroid-V2 across different LLMs. The results presented in Figure 5 show the comparison of success rate and reversed redundancy ratio between AutoDroid-V2 and the baseline models for different LLMs: Llama3.2-3B, Qwen2.5-7B, and Llama3.1-8B, where AutoDroid-V2 achieves 44.6%, 50.0%, 54.4% success rate respectively. The reversed redundancy ratio of AutoDroid-V2 is stable, ranging from 90.5% to 93.0%. This is due to the generated scripts being designed to be straightforward and focused on task completion without including exploration or trial-and-error steps.
@@ 5. Discussion
Another possible limitation of AutoDroid-V2 is the difficulty of handling highly dynamic user interfaces, such as those in web search engines and games. In these apps, it is difficult to plan the actions in advance, since the future states are usually unpredictable. In concept, AutoDroid-V2 can deal with such dynamic user interfaces, but it may need to regenerate the plan for each step, leading to degraded efficiency. Our future work will focus on how to seamlessly integrate the step-wise reasoning approach with our script-based approach. Future agents are expected to flexibly switch between the two modes for different types of user interfaces, handling dynamic content effectively while maintaining efficiency.
@@ 6. Conclusion
Acknowledgement
This work is supported by the National Key R&D Program of China (No.2022YFF0604500), National Natural Science Foundation of China (Grant No.62272261), Xiaomi Foundation, Wuxi Research Institute of Applied Technologies, Tsinghua University (Grant No.20242001120), and a grant from the Guoqiang Institute, Tsinghua University.